LBTO Techincal/D/Service Observing Night

OSA: D.G.Huerta

ISp: J.Morris

Manager: J.Urban

ISA: B.Rothberg, D.Thompson, J.Hill (Tucson)

================================================

UT 11:26 (Nov20) - scope configured for LUCI-2 at start

UT 12:48 - updated PCS with MODS2 0.6 degree rotation correction

UT 01:10 - moving to HD210501 - test of LUCI-2 throughput

UT 01:18 - doing a pointing check as guidestar acquisition failed twice

UT 01:19 - resending preset to HD210501 again

UT 01:40 - Error - voltage didn't stabilize at 2.2 micron, script aborted. - last image was 0012

UT 01:43 - Aborted, clicked skip to the 2.2micron @K 210 and started - it did an extra -20" dither

UT 01:44 - Aborted with voltage error again

UT 01:44 - manually putting in 2.2micron

UT 01:45 - error committing instrument setup with 2.2 micron @210 grating

UT 01:47 - again, another error commiting instrument setup - trying to set to 1.65 then 2.21, then 2.20

UT 01:49 -didn't like 2.21, trying 2.19

UT 01:50 - initializing grating, trying 2.20 - failed again

UT 01:52 - trying 2.22 - failed again

UT 01:55 - 2.3 worked, trying 2.25

UT 01:55 - 2.25 didn't work, trying 2.29, worked, trying 2.26

UT 01:57 - 2.26 failed, trying 2.27, worked

UT 01:58 - trying 2.24, failed trying again

UT 02:01 - i'm continuing script at 2.27 microns - note that everything is -20" off from where it should be in script - and we skipped one dither position in the K observing

UT 02:25 - Failed to set grating 210 to 2.2microns

UT 02:31 - still can't get anything at < 2.27 microns to stabilize

UT 02:33 - aborting script - reloading version with 2.27 micron for field stop, and redoing field stop part of script

UT 02:34 - sending preset again - going to line up again on LS2" slit, then skip in script down to field stop section and redo it

UT 02:47 - restarted N30 field stop portion

UT 03:03 - LUCI2 testing completed for throughput

========================================

UT 03:04 - LUCI-2 imaging of M31 - RMWagner

UT 03:33 - Completed M31 imaging field - moving to sky position

UT 03:34 - missed the guidestar - aborting script

UT 03:36 - resending M31-sky script

UT 04:03 - Script finished - some variable seeing towards the end

========================================

UT 04:07 - Moving to Stone A3 field for LUCI2 transform tests

UT 04:10 - N3.75 Camera, Pointing Mask, Br Gamma filter

UT 04:33 - some issues in getting transform running - seeing still highly variable

UT 04:37 - starting again

UT 04:39 - #0069 blank (blinds in)

UT 04:43 - starting again

UT 04:48 - 0070 first image

GS 8 = 0072

GS 2 = 0073 (binary?)

GS 5 = 0074

GS20 = star too far out

UT 05:01 - relodaing with start=4

GS 24 = stopped again, too far of a move

UT 05:07 - restarting again, start=3

GS 10 = 0075

GS 11 = 0076

GS 9 = 0077

GS 8 = stopped on this too

UT 05:17 - restarting again, start=7 - clouds visible across southern sky

GS 5 = 0078

GS 10 = 0079

GS 11 = 0080

GS9 = No guidestar found in box - could be too faint

UT 05:26 - restarting again, start=11

GS 4 = err = Resmsg[2] = error: isPosAccessible: couldn't move probe to position (out of patrol field) (SFP) (x,y): 134.88, -26.3228 mm.

UT 05:29 - restarting again, start=12

GS 0 = 0081

UT 05:39 - moving to StoneD to do similar testing with field we've done before - using Stone_D_new_Oct27.list

GS0 = 0082

offset 30 = 0083

GS4 = 0084 PA=0

GS12 = 0085

GS23 = 0086

GS13 = 0087

GS22 = 0088

GS5 = 0089

GS2 = 0090

GS11 = 0091

GS29 = 0092

GS24 = 0093

GS30 = 0094

GS25 = 0095

GS4 = 0096 PA=15

GS12 = 0097

GS23 = 0098 --- seeing 1" or less

GS13 = 0099

GS22 = 0100

GS17 = 0101

GS5 = 0102

GS11 = 0103

GS29 = 0104

GS24 = 0105

GS30 = 0106 -seeing bloom

GS4 = 0107 PA=30

GS12 = 0108

GS23 = 0109

GS13 = 0110

GS22 = 0111

GS5 = 0112

GS11 = 0113

GS29 = 0114

GS24 = 0115

GS30 = failed preset - this happened Oct 27 too, so its ok

================================================

UT 06:56 - switching to MODS2 - putting LUCI2 in safe config

UT 07:13 - mods2r.201121.0002 - test exposure

UT 07:15 - put in ND1.5 filter - 1sec ~9300 counts

UT 07:18 - 4 sec ~22k peak

UT 07:21 - D.Miller remote test image -0005

UT 07:22 - starting collection

UT 07:27 - 006,007 guideprobe in way

UT 07:35 - exited on GS12, restarted with adjustments made to moving guideprobe

UT 07:43 - GS23 failed

UT 07:44 - restarting with start=3

UT 07:52 - changed exposure time to 3seconds to prevent saturation - GS5 Pa=180 is last 4sec exposure

UT 07:57 - failed on GS29 PA=180, restarting with restart=9

UT 08:00 - failed on GS24, restart=10

UT 08:02 - failed on GS30, restart=11

UT 08:04 - failed on GS25,restart=12

UT 08:07 - seeing subarcsecond

UT 08:08 - failed on GS12 PA=195 - changing exptime to 2sec

UT 08:10 - failed on GS23, restart=15

UT 08:22 -failed on GS29 PA=195, restart=21

UT 08:26 - failed on GS24 PA=195, restart=22

UT 08:28 - failed on GS30 PA=195, restart=23

UT 08:34 - failed on GS23 PA=195, restart = 26

UT 08:40 - seeing degraded a bit

UT 08:44 - failed on GS11 PA=210, restart=30

UT 08:50 - having trouble collimating - got an error after 7 iteratoins - continuing

UT 08:52 - failed on GS29 PA=210, restart=31

UT 08:55 - failed on GS24 PA=210, restart=32

UT 09:00 - failed on GS30 and GS20 PA=210

=================================

UT 09:01 - moving to V959Mon (LST 05:30 script) - science field for RMWagner-OSURC

UT 09:04 - seeing 1" in guider

UT 09:11 - not clear which is the actual target - is it the bright star I centered on, or the slightly fainter blob to bottom left of it...

UT 09:15 - exposure started - will know if its correct target after first spectra reads out

UT 09:29 - i'm assuming this is it as I see bright emission lines

UT 09:50 - completed target - moving to G191B2b spec-phot std

=================================

UT 09:59 - centering up g191b2b a bit in slit +-0.6" x

UT 10:00 - starting g191b2b spectra

UT 10:11 - finished std

==================================

UT 10:12 - moving to V380 Orionis - Rothberg target

UT 10:23 - need to take a shorter exp field image

UT 10:26 - starting spectra (r:0044-46, b:0008-0010)

UT 10:32 - had three IMCS red lock failures, elevation = 44

UT 10:38 - again no lock-on for red

UT 10:47 - ran through the procedure to power down and up lasers manually, still no joy, red imcs will not lock

UT 11:09 - went back to G191b2b to test, worked ok, came back to V380, acq worked ok, now trying spectroscopy again

UT 11:13 - MODS2 still having issues with RED IMCS lock, switching back to LUCI2 for end of night

UT 11:15 - IT 5767 for MODS2 Red IMCS lock

UT 11:33 - moving to WAS49

UT 11:48 - acq complete - taking K, then H G210 spectra

UT 12:00 - noticed flexure compensation appears to be off in GUI, but script says its on - header says Flexure comp is off - appears to have turned off after acquisition completed (on at 119, off 120-125)

UT 12:05 - flexure compensation went back ON after switch to H-band

UT 12:10 - bottom half of script vanished after I had put in a pause and removed it.

UT 12:14 - guiding paused on -60" offset, but I had already done this offset earlier - I suspect the "correction" when script vanished, might have done two dither moves instead of one

UT 12:22 - going to run through script again, manually adding back in +60" +X dither to get back on track

UT 12:23 - re-running K, flexure compensation turned itself off

UT 12:35 - removed a pause and script vanished again and can't continue. -Note flexure compensation is back on for last two K-G210 exposures in sequence (136-137)

UT 12:39 - 0.5-0.7" in guider

UT 12:57 - Moving to telluric

UT 01:16 - finished, closing enclosure then will do cals

UT 01:23 - LUCI2 - cals 0.5" slit K

UT 01:27 - commit error for cal setup - had to restart scheduler

UT 01:28- running slitflats 0.8" MODS2

UT 01:44 - LUCI2 instrument manager frozen?

UT 01:48 - Called Dave T. - unable to restart software or do anything, however GEIRS works

UT 01:49 - Running slitless pixel flats MODS2

UT 02:02 - Dave T. restarted instrument manager -all seems ok - running 0.5" G210 @K2.27 cals

UT 02:09 - MODS2 arc lamps

UT 02:19 - MODS2 8k biases

UT 02:27 - starting LUCI2 0.5"G210H cals

UT 02:28 - putting mods2 to sleep

UT 02:36 - doing h flats and just the argon arcs

UT 02:41 - putting LUCI2 into safe config, N375, blindsx2, mirror, flex off, cal unit out, leaving 0.5" slit in FPU

-- %USERSIG{BarryRothberg - 2015-11-20}%

Comments

 
Topic revision: r6 - 21 Nov 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback